I have in a tank 200 liters of water, I'm pouring 3 ounces of anti foam liquid. I want to know how many parts per million of anti foam liquid is in the water.
1
Expert's answer
2017-05-29T09:48:10-0400
To find parts per million (ppm) we have to divide mass of solute by mass of solution and then multiply by one million:
ppm = (mass of solute / mass of solution) * 1 000 000
Note that all masses have to be expressed in the same units. Actually it doesn’t matter what units to use, but we will use kilograms as it is generally accepted in science.

1 ounce = 0.028 kg. So the mass of anti foam agent is 3 * 0.028 = 0.084 kg.

Mass of water = volume of water * density of water. The density of water is about 1 kg/L. Than the mass of water in a tank is 200 L * 1 kg/L = 200 kg.

Mass of solution = mass of water + mass of anti foam liquid = 200 kg + 0.084 kg = 200.084 kg.
Actually in this case we can neglect the mass of anti foam because it is extremely small comparing to the mass of water, but I just want to show all steps properly.

Now, calculate ppm:

Ppm = (0.084 kg / 200.084 kg) * 1000000 = 420.


Answer:
There are 420 parts per million of anti foam liquid in the water.